OpenStudy (anonymous):

Fill in the missing numbers to form equivalent rates. 2/30 = __/15 = __/45

OpenStudy (anonymous):

hi sam, so this is a basic ratio problem - do you know how to start it?

OpenStudy (anonymous):

no can you gimme the answer

OpenStudy (anonymous):

Sorry sam, we're here to help you understand the problem, not give you the answer. Let me start it for you though:

OpenStudy (anonymous):

\[2/30 = 1/15\] The denominator (15) is half of the original (30) so the numerator (1) needs to be also half. Can you figure out the other one?

OpenStudy (anonymous):

no i cant

OpenStudy (anonymous):

Got a guess? If 2/30 = 1/15 then X/45 - got a guess for X?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

i get it

OpenStudy (anonymous):

Close! You're almost there - 2/30 = 1/15 right? So now, if I need something over 45 to equal that...

OpenStudy (anonymous):

What answer did you come up with?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

No, sorry - 4/60 = 2/30

OpenStudy (anonymous):

So knowing that, now you've got 4/60 = X/45 = 2/30 = 1/15... Leaving what number for X?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Yes! Well done!

OpenStudy (anonymous):

what one

OpenStudy (anonymous):

It's three - to check if you divide both sides of the fraction 3/45 by three you get 1/15

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

You're welcome.